# EMBRACE: Shaping Inclusive Opinion Representation by Aligning Implicit Conversations with Social Norms

### v1

@@ 4.1 Evaluating Normative Alignment in Implicatures Conversations
We demonstrate the interplay between human and LLM responses in various scenarios to compare the distinct behavior of assistance stance between implicit and explicit opinion as shown in Figure 2. In general, humans tend to follow the normative expectation of disagreeing with toxic content, especially when discourse is explicit. In particular Expert humans show high disagreement rates toward explicit opinion, reflecting a stronger normative alignment. Interestingly, LLMs have a higher likelihood of agreement when conversation is explicitly has harmful opinion, potentially due to surface-level alignment. In contrast, responses to implicit discourse elicit more neutral stances from LLMs, suggesting hesitation or ambiguity in detecting subtler expressions. All comparisons are statistically significant based on a chi-square test () as shown in Appendix B. Moreover, we analyze the confidence markers associated with the assistance responses. As shown in figure 3, LLMs tend to respond more cautiously, using “Refuse to Engage” or neutral tones more often, and expressing confidence more explicitly. In contrast, humans disagree openly in implicit contexts but rarely tag their confidence. Overall, the vast majority of human certainty is marked as None (91.1%), indicating that humans do not explicitly express confidence as often.
@@
As shown in figure 3, LLMs tend to respond more cautiously, using “Refuse to Engage” or neutral tones more often, and expressing confidence more explicitly. In contrast, humans disagree openly in implicit contexts but rarely tag their confidence. Overall, the vast majority of human certainty is marked as None (91.1%), indicating that humans do not explicitly express confidence as often.
Stance transitions in implicit conversations
We analyze the turning point of stance within the conversation as shown in Figure 4. Mainly, it illustrates the distribution of user stance positions within conversations involving human and LLM-generated responses. The y-axis represents the normalized position of each user’s turn, with higher values indicating later turns. Across both assistant types, agree and neutral stances are expressed in later parts of the conversation. However, two key patterns can be noticed. First, initial stances in human dialogues occur significantly earlier when users hold implicit opinions, indicating an early assertion of viewpoint under ambiguity. Second, users are more likely to express disagree stances earlier in conversations with humans than with LLMs, especially when opinions are implicit. For LLMs, the only significant shift appears in the neutral stance, where users with implicit opinions tend to reach neutrality earlier. These patterns suggest users exhibit greater conversational assertiveness, either through disagreement or early opinion assertion when responding to human assistants, while interactions with LLMs shows more delayed or neutral positioning. We validated the significance of our comparison and conducted Mann-Whitney U tests comparing the relative timing of user stances between explicit and implicit opinion contexts (Appendix B).
@@ 4.2 Model Performance across Implicit Training Portions
Figure 5 complements these results by showing that both Mistral and Linear models maintain low false positive rates (FPR), especially beyond 10% inclusion. Notably, MLP models exhibit a sharp spike in FPR at 0% and zero-shot settings, underscoring their inability to generalize without the inclusion of implicit cues. This overprediction of the Agree class in norm-sensitive contexts demonstrate poor calibration and indicates risk of norm violation. In contrast, LLaMA3 maintains a low FPR at these early settings, but this is linked with low macro F1 scores (see Table 2), suggesting underprediction or overly conservative behavior rather than calibrated learning, which is a different type of failure mode. When comparing the implicit and explicit False Positive Rates (FPRs) within each model, both PU-learned models (Linear and MLP) and LLaMA3 exhibit a relatively small FPR gap across different discourse styles. This indicates that these models behave consistently, regardless of whether the language used is subtle or overt. In contrast, the Mistral model exhibits a larger FPR gap, especially at low inclusion levels, which suggests a bias toward surface-level (explicit) cues.
The narrower FPR disparity seen in the PU models and LLaMA3 indicates better fairness and robustness in adapting to stylistic variations. This low FPR is due to different reasons, as PU models benefit from fairness-aware training of implicit and explicit groups. While LLaMA3’s uniform behavior of FPR between explicit and implicit opinions suggests that the model tends to adopt a conservative stance by avoiding agreement even when it may be the correct stance.
@@ 5 Discussion and Implications
As illustrated in Figures 2 and 3, assistant stance response behaviors differ across implicit and explicit user opinions. Figure 2 reveals that LLMs have a higher rate of agreement with explicit extreme toxic opinion, compared to implicit toxicity. While, expert assistants humans show more stable disagreement regardless of implicit or explicit misaligned norms. By zooming in on the neutral user opinion as shown in Figure 7, Appendix A, human assistants are more likely to disagree, while LLMs tend to still be agreeable. This confirm our experiment design to focus on extreme clear cases of implicit toxic opinion to facilitate the overall examination of stance as a means to evaluate the social norms. This behavior of complicity in LLMs’ responses, even toxic opinions, has been underscored in previous studies as “sycophancy” (Hong et al., 2025; Cheng et al., 2025; Rrv et al., 2024), where LLMs show agreeable behavior with users’ statements. However, our findings extend this line of work by examining agreement in the presence of implicit opinion cues, such as implicatures and indirect expressions of norm misaligned context (toxicity). Unlike prior studies that focus on overt stance shifts, we demonstrate that LLMs remain agreeable even in subtly toxic or norm mismatch contexts, particularly when opinions are implicitly framed. Complementing this, Figure 3 shows that in implicit opinions, human assistants tend to respond to initial or neutral stances with clear disagreement and high confidence, whereas LLMs often either refuse to engage or express uncertainty. These trends underscore a normative alignment gap in LLM responses, where human assistants tend to maintain a more decisive and oppositional stance toward problematic content that is toxic. At the same time, LLMs display an inconsistent stance of neutrality and usually tend to agree with those toxic misaligned norms and signaling and lower confidence when facing implicit toxicity. In contrast to prior work that analyzes confidence markers in isolation (Röttger et al., 2024), our analysis reveals that focusing solely on refusal or uncertainty overlooks how models may simultaneously express stance alignment, espicialy the neutral or agreeable stances toward norm violating content. This behaviour sheds light on a fixed or superficial reliance on neutral responses, which might not be a sufficient safeguard in value sensitive conversations, especially when toxicity or bias is embedded through implicature or indirect opinion expression. Our findings advocate for integrating stance analysis with confidence calibration to better evaluate normative alignment in implicit contexts.
@@
Dynamics of user stance within conversation narratives Building on our examination of user stance toward assistant replies, we further analyze the user reaction towards the assistant. As shown in figure 6, it can be seen that when extreme opinion is overt, users recognize the assistant’s corrective or balanced stance and respond supportively. On the other hand, implicit extreme opinion has greater user disagreement, potentially because the harm is debatable (Especially toward human assistants). Unlike human assistants, LLMs provoke more user disagreement when responding to explicit extreme opinions, particularly when they remain neutral. This behavior can be explained through the “Elaboration Likelihood Model (ELM) of persuasion” (Petty and Cacioppo, 1986), which states that attitudes change occurs through either central (deep) or peripheral (surface-level) processing route. In the case of LLM generated responses, users may fail to engage in deep processing if the assistant’s message lacks perceived credibility or personal relevance. Instead, users usually rely on peripheral cues (surface-level), such as tone of confidence or refusal to engage in the conversation, as we observed in Figure 3, where the LLM frequently adopts a refusal tone. Consequently, users are less likely to shift their stance or revise their disagreement in response to the LLM, unless it presents highly credible or reasoned arguments. This tendency has been supported in different ways by surface-level stance interactions (Aldayel and Magdy, 2022) or as demonstrated by (Gallegos et al., 2025) through user perspective on labeled AI responces.
@@
A closer examination of the results in Table 2 and Figure 5 shows that scaling the inclusion of implicit conversational data results in measurable improvements in both performance and calibration of FPRs. As can be seen, Mistral’s overall performance is enhanced compared to zero-shot and remains robust at partial training levels, achieving high F1 scores (above 0.93 from as low as 10% inclusion) while maintaining a low False Positive Rate (FPR), which is an indicator of reliability in norm sensitive classification. Also, linear PU models trained on implicit opinions has consistently low FPRs demonstrating the benefit of even shallow architectures (linear) when trained on implicit supervision. In contrast, deeper MLP models remain less reliable, with high FPRs, suggesting that more complex models may require additional regularization or architectural adjustments to handle implicit nuance effectively (Topic level discrimination). The 30% inclusion portion is a critical threshold as below that threshold, models struggle with implicit opinion patterns. While above 30%, Mistral and Linear models show consistent model behavior. We further verify the performance between models comparisons using MacNamer’s test in Tables 9, 10 (Appendix D) which support this behaviour with evidence of reduced overgeneralization errors past this point. Thus, a monitored and balanced inclusion of implicit data improves accuracy and minimizes false agreement with harmful perspectives.
@@ 6 Conclusion
The datasets used such as DialogConan, ToxicChat may reflect sociocultural norms that are specific to certain communities or platforms. Thus, the generalizability of the normative alignment framework across diverse cultural and linguistic contexts remains limited and needs to be considered in future cross-cultural studies. Additionally, our social norm evaluation used a few LLMs (Mistral, LLaMA3) transformer-based models that tend to memorize factual and normative patterns from their extensive pretraining. However, our current analysis scope does not empirically assess how temporal aspects of model training, or the evolving nature of norms within training data, might impact their alignment with socially expected stances. As future direction need to consider examining this aspect, especially the temporal shifts in normative behavior and their impact on stance consistency.
@@
This study aims to advance equitable inclusive of opinion representation in conversational models by including implicit opinion and using stance as means to evaluate normative alignment. Motivated by ethical computing principles such as ACM Code of Ethics Principle 1.4 (“Be fair and take action not to discriminate”), this study seeks to evaluate the implicit language through which conversational models may reinforce norm-violating or harmful views. Although, we recognize that any biased or poorly designed community language models can unintentionally reinforce stereotypes. We highlight that our framework does not view human disagreement as mere noise but as an important reflection of social norms. Our objective is to encourage value-sensitive, inclusive design without silencing diverse yet respectful viewpoints.
@@ Appendix A Task Formation
In line with the EMPRACE framework’s emphasis on equitable inclusion of implicit opinion expressions, we design an annotation task to explore the boundaries of extreme opinions, particularly those conveyed through toxic language. We treat toxicity (explicit and implicit) as a heightened form of stance expression. We ground the annotation logic in a normative framework, where toxic content (such as extreme ideological disagreement) is expected to be opposed in healthy discourse. In the light of (Grice, 1975), which references Implicit Attitude Theory, this annotation specification aims to better evaluate subtle language patterns as meaningful indicators of user and follow-up assistant stance, rather than dismissing them as noise. By zooming in on cases where the user stance is neutral (Figure 7), we observe a noticeable divergence between human and LLM assistant responses as humans are more likely to adopt a disagreeing stance, whereas LLMs disproportionately favor agreement or neutrality. This can be further illustrated with in human-human interaction as shown in Figure 7(b), as with in Conan (Expert human) assistant, the rate of disagreement from these experts is higher, in comparison with open conversations tweetscontext data, this might draw on the nature of the data, as experts might be accustomed to expect the worse intention and fight back in the conversations. Thus, our annotation schema is designed to represent these nuanced aspects and relate them with a normative stance expectation: toxicity, in all its forms, is presumed to warrant disagreement, allowing us to trace how language models or humans respond to extremity across social contexts.
@@ A.1 Annotation Process
For user responses, the annotation included two dimensions. First, Toxicity was categorized as Explicit Toxicity, Implicit Toxicity, or Neutral, to capture both overt and subtle harms. Second, Stance toward the assistant’s proposition was labeled as Agree and Support, Disagree and Oppose, Elaborate or Neutral, Initial Message, or Start a New Topic. This multi-layered annotation process enables fine-grained modeling of toxicity as implicit opinion expression and normative alignment. Further details of the guidelines and data can be accessed through (anonymized link111anonymized link for peer Reviewhttps://osf.io/2azn5/files/osfstorage?view_only=bd8f756bb2e849a1b5102953cf33a775). We used Labelbox to run the annotation process and crowd-sourcing. Each turn has been annotated by two annotators recruited via Labelbox222https://labelbox.com/services/annotation-services. Labelbox provides a managed workforce of trained annotators in 40-hour blocks at a rate of $8 USD per hour. Each annotator must pass benchmark tests and quality assurance checks before contributing to the project and a third reviewer (one of the authors) to verify annotation consistency and ensure pipeline quality.
Based on the Kappa scores reported in Table 3, most agreement levels fall within the moderate (0.41–0.60) to substantial (0.61–0.80) range (McHugh, 2012). Assistant confidence and stance annotations show substantial and moderate agreement, respectively, for assistant-human conversations ( and ), and moderate agreement for assistant-chat data ( and ). For user-based annotations, user-chat data exhibited substantial agreement on user stance () and moderate agreement on implicit opinion (), with the highest average Kappa overall (). In contrast, user-human data showed only fair agreement on implicit opinion () and moderate agreement on user stance (), with the lowest overall average agreement ().

### v2

@@ 4.1 Evaluating Normative Alignment in Implicatures Conversations
We demonstrate the interplay between human and LLM responses in various scenarios to compare the distinct behavior of assistance stance between implicit and explicit opinion as shown in Figure 2. In general, humans tend to follow the normative expectation of disagreeing with toxic content, especially when discourse is explicit. In particular, Expert humans show high disagreement rates toward explicit opinion, reflecting a stronger normative alignment. Interestingly, LLMs have a higher likelihood of agreement when the conversation explicitly has a harmful opinion, potentially due to surface-level alignment. In contrast, responses to implicit discourse elicit more neutral stances from LLMs, suggesting hesitation or ambiguity in detecting subtler expressions. All comparisons are statistically significant based on a chi-square test () as shown in Appendix B. To verify that these patterns are not artifacts of mixed datasets, we analyzed the two LLM subsets separately (Figure 3).Across both WildChat (ChatGPT) and ToxicChat (Vicuna-7B), Figure 3 replicates the pattern in Figure 2. Particularly, LLMs show higher agreement with explicit harmful opinions and more neutral responses to implicit ones, confirming surface-level rather than contextual norm alignment. Across models, ChatGPT (WildChat) exhibits slightly higher neutrality, whereas Vicuna-7B (ToxicChat) shows stronger disagreement with explicit opinions. Moreover, we analyze the certainty markers associated with the assistance responses. As shown in figure 4, LLMs tend to respond more cautiously, using “Refuse to Engage” or neutral tones more often, and expressing certainty more explicitly. In contrast, humans disagree openly in implicit contexts but rarely tag their certainty. Overall, the vast majority of human certainty is marked as None (91.1%), indicating that humans do not explicitly express certainty as often.
@@
As shown in Figure 4, LLMs tend to respond more cautiously, using “Refuse to Engage” or neutral tones more often, and expressing certainty more explicitly. In contrast, humans disagree openly in implicit contexts but rarely tag their certainty. Overall, the vast majority of human certainty is marked as None (91.1%), indicating that humans do not explicitly express certainty as often.
@@
Stance transitions in implicit conversations
We analyze the turning point of stance within the conversation as shown in Figure 5. Mainly, it illustrates the distribution of user stance positions within conversations involving human and LLM-generated responses. The y-axis represents the normalized position of each user’s turn, with higher values indicating later turns. Across both assistant types, agree and neutral stances are expressed in later parts of the conversation. However, two key patterns can be noticed. First, initial stances in human dialogues occur significantly earlier when users hold implicit opinions, indicating an early assertion of viewpoint in the face of ambiguity. Second, users are more likely to express disagree stances earlier in conversations with humans than with LLMs, especially when opinions are implicit. For LLMs, the only significant shift appears in the neutral stance, where users with implicit opinions tend to reach neutrality earlier. These patterns suggest users exhibit greater conversational assertiveness, either through disagreement or early opinion assertion when responding to human assistants, while interactions with LLMs show more delayed or neutral positioning. We validated the significance of our comparison and conducted Mann-Whitney U tests comparing the relative timing of user stances between explicit and implicit opinion contexts (Appendix B).
@@ 4.2 Model Performance across Implicit Training Portions
Figure 6 complements these results by showing that both Mistral and Linear models maintain low false positive rates (FPR), especially beyond 10% inclusion. Notably, MLP models exhibit a sharp spike in FPR at 0% and zero-shot settings, underscoring their inability to generalize without the inclusion of implicit cues. This overprediction of the Agree class in norm-sensitive contexts demonstrates poor calibration and indicates risk of norm violation. In contrast, LLaMA3 maintains a low FPR at these early settings, but this is linked with low macro F1 scores (see Table 2), suggesting underprediction or overly conservative behavior rather than calibrated learning, which is a different type of failure mode. When comparing the implicit and explicit False Positive Rates (FPRs) within each model, both PU-learned models (Linear and MLP) and LLaMA3 exhibit a relatively small FPR gap across different discourse styles. This indicates that these models behave consistently, regardless of whether the language used is subtle or overt. In contrast, the Mistral model exhibits a larger FPR gap, especially at low inclusion levels, which suggests a bias toward surface-level (explicit) cues.
The narrower False Positive Rate gap observed in PU models and LLaMA-3 indicates more balanced treatment of implicit and explicit language, suggesting higher fairness consistency. This low FPR is due to different reasons, as PU models benefit from fairness-aware training of implicit and explicit groups. While LLaMA3’s uniform behavior of FPR between explicit and implicit opinions suggests that the model tends to adopt a conservative stance by avoiding agreement even when it may be the correct stance.
@@ 5 Discussion and Implications
Our findings reveal that LLMs exhibit surface-level alignment by agreeing with explicitly toxic opinions. At the same time, humans maintain stable normative disagreement. This indicates that model alignment remains superficial when norms are implied rather than overtly stated. As illustrated in Figures 2 and 4, assistant stance response behaviors differ across implicit and explicit user opinions. Figure 2 reveals that LLMs have a higher rate of agreement with explicit extreme toxic opinion, compared to implicit toxicity. While, expert assistants humans show more stable disagreement regardless of implicit or explicit misaligned norms. By zooming in on the neutral user opinion as shown in Figure 8, Appendix A, human assistants are more likely to disagree, while LLMs tend to still be agreeable. This confirm our experiment design to focus on extreme clear cases of implicit toxic opinion to facilitate the overall examination of stance as a means to evaluate the social norms. This behavior of complicity in LLMs’ responses, even toxic opinions, has been underscored in previous studies as “sycophancy” (Hong2025-pz; Cheng2025-wk; Rrv2024-bq), where LLMs show agreeable behavior with users’ statements. However, our findings extend this line of work by examining agreement in the presence of implicit opinion cues, such as implicatures and indirect expressions of norm misaligned context (toxicity). Unlike prior studies that focus on overt stance shifts, we demonstrate that LLMs remain agreeable even in subtly toxic or norm mismatch contexts, particularly when opinions are implicitly framed. Complementing this, Figure 4 shows that in implicit opinions, human assistants tend to respond to initial or neutral stances with clear disagreement and high certainty, whereas LLMs often either refuse to engage or express uncertainty. These trends underscore a normative alignment gap in LLM responses, where human assistants tend to maintain a more decisive and oppositional stance toward problematic content that is toxic. At the same time, LLMs display an inconsistent stance of neutrality and usually tend to agree with those toxic misaligned norms and signaling, and lower certainty when facing implicit toxicity. In contrast to prior work that analyzes certainty markers in isolation (Rottger2024-vv), our analysis reveals that focusing solely on refusal or uncertainty overlooks how models may simultaneously express stance alignment, espicially the neutral or agreeable stances toward norm violating content. This behaviour sheds light on a fixed or superficial reliance on neutral responses, which might not be a sufficient safeguard in value sensitive conversations, especially when toxicity or bias is embedded through implicature or indirect opinion expression. Our findings advocate for integrating stance analysis with certainty calibration to better evaluate normative alignment in implicit contexts.
@@
Dynamics of user stance within conversation narratives Building on our examination of user stance toward assistant replies, we further analyze the user reaction towards the assistant. As shown in figure 7, it can be seen that when extreme opinion is overt, users recognize the assistant’s corrective or balanced stance and respond supportively. On the other hand, implicit extreme opinion has greater user disagreement, potentially because the harm is debatable (Especially toward human assistants). Unlike human assistants, LLMs provoke more user disagreement when responding to explicit extreme opinions, particularly when they remain neutral. This behavior can be explained through the “Elaboration Likelihood Model (ELM) of persuasion” (Petty1986-ub), which states that attitude change occurs through either central (deep) or peripheral (surface-level) processing route. In the case of LLM generated responses, users may fail to engage in deep processing if the assistant’s message lacks perceived credibility or personal relevance. Instead, users usually rely on peripheral cues (surface-level), such as tone of certainty or refusal to engage in the conversation, as we observed in Figure 4, where the LLM frequently adopts a refusal tone. Consequently, users are less likely to shift their stance or revise their disagreement in response to the LLM, unless it presents highly credible or reasoned arguments. This tendency has been supported in various ways by surface-level stance interactions (Aldayel2022-ew) or as demonstrated by (Gallegos2025-zz) through a user’s perspective on labeled AI responses.
@@
A closer examination of the results in Table 2 and Figure 6 shows that scaling the inclusion of implicit conversational data results in measurable improvements in both performance and calibration of FPRs. As can be seen, Mistral’s overall performance is enhanced compared to zero-shot and remains robust at partial training levels, achieving high F1 scores (above 0.93 from as low as 10% inclusion) while maintaining a low False Positive Rate (FPR), which is an indicator of reliability in norm sensitive classification. Additionally, linear PU models trained on implicit opinions consistently exhibit low FPRs, demonstrating the benefits of even shallow architectures (linear) when trained on implicit supervision. In contrast, deeper MLP models remain less reliable, with high FPRs, suggesting that more complex models may require additional regularization or architectural adjustments to handle implicit nuance effectively (topic-level discrimination). The 30% inclusion portion is a critical threshold as below that threshold, models struggle with implicit opinion patterns. While above 30%, Mistral and Linear models show consistent model behavior. We further verify the performance between models comparisons using MacNamer’s test in Tables 9, 10 (Appendix D) which support this behaviour with evidence of reduced overgeneralization errors past this point. Thus, a monitored and balanced inclusion of implicit data improves accuracy and minimizes false agreement with harmful perspectives.
@@ 6 Conclusion
The datasets used such as DialogConan, ToxicChat may reflect sociocultural norms that are specific to certain communities or platforms. Thus, the generalizability of the normative alignment framework across diverse cultural and linguistic contexts remains limited and needs to be considered in future cross-cultural studies. Additionally, our social norm evaluation used a few LLMs (Mistral, LLaMA3) transformer-based models that tend to memorize factual and normative patterns from their extensive pretraining. While human and LLM dialogues originate from distinct contexts, this design intentionally preserves ecological validity, reflecting how implicit opinions naturally arise in real-world conversations. Future work will extend EMBRACE to controlled settings using identical prompts for both humans and LLMs to strengthen cross-domain comparability. However, our current analysis scope does not empirically assess how temporal aspects of model training, or the evolving nature of norms within training data, might impact their alignment with socially expected stances. As a future direction, it would be beneficial to consider examining this aspect, particularly the temporal shifts in normative behavior and their impact on stance consistency.
@@
This study aims to advance equitable inclusive of opinion representation in conversational models by including implicit opinion and using stance as means to evaluate normative alignment. Motivated by ethical computing principles such as ACM Code of Ethics Principle 1.4 (“Be fair and take action not to discriminate”), this study seeks to evaluate the implicit language through which conversational models may reinforce norm-violating or harmful views. Although we recognize that any biased or poorly designed community language models can unintentionally reinforce stereotypes. We emphasize that our framework does not view human disagreement as mere noise, but rather as an important reflection of social norms. Our objective is to encourage value-sensitive, inclusive design without silencing diverse yet respectful viewpoints. Annotators were briefed on handling toxic material and provided with opt-out options.
| Data | Avg. | Kappa | Kappa |
| Kappa | Conf. Asst. | Stance Asst. | |
| Assistant_LLMchats | 0.571 | 0.61 | 0.53 |
| Assistant_human | 0.569 | 0.6326 | 0.506 |
| Data | Avg. | Kappa | Kappa |
| Kappa | Imp. Op. | Stance User | |
| User_LLMchat | 0.7 | 0.57 | 0.83 |
| User_human | 0.475 | 0.40 | 0.55 |
@@ Appendix A Task Formation
In line with the EMBRACE framework’s emphasis on equitable inclusion of implicit opinion expressions, we design an annotation task to explore the boundaries of extreme opinions, particularly those conveyed through toxic language. We treat toxicity (explicit and implicit) as a heightened form of stance expression. We ground the annotation logic in a normative framework, where toxic content (such as extreme ideological disagreement) is expected to be opposed in healthy discourse. In the light of (grice1975logic), which references Implicit Attitude Theory, this annotation specification aims to better evaluate subtle language patterns as meaningful indicators of user and follow-up assistant stance, rather than dismissing them as noise. By zooming in on cases where the user stance is neutral (Figure 8), we observe a noticeable divergence between human and LLM assistant responses as humans are more likely to adopt a disagreeing stance, whereas LLMs disproportionately favor agreement or neutrality. This can be further illustrated with in human-human interaction as shown in Figure 8(b), as with in Conan (Expert human) assistant, the rate of disagreement from these experts is higher, in comparison with open conversations tweetscontext data, this might draw on the nature of the data, as experts might be accustomed to expect the worse intention and fight back in the conversations. Thus, our annotation schema is designed to represent these nuanced aspects and relate them with a normative stance expectation: toxicity, in all its forms, is presumed to warrant disagreement, allowing us to trace how language models or humans respond to extremity across social contexts.
@@ A.1 Annotation Process
For user responses, the annotation included two dimensions. First, Toxicity was categorized as Explicit Toxicity, Implicit Toxicity, or Neutral, to capture both overt and subtle harms. Second, Stance toward the assistant’s proposition was labeled as Agree and Support, Disagree and Oppose, Elaborate or Neutral, Initial Message, or Start a New Topic. This multi-layered annotation process enables fine-grained modeling of toxicity as implicit opinion expression and normative alignment. Further details of the guidelines and data can be accessed through (anonymized link222anonymized link for peer Reviewhttps://osf.io/2azn5/files/osfstorage?view_only=bd8f756bb2e849a1b5102953cf33a775). We used Labelbox to run the annotation process and crowd-sourcing. Each turn has been annotated by two annotators recruited via Labelbox333https://labelbox.com/services/annotation-services. Labelbox provides a managed workforce of trained annotators in 40-hour blocks at a rate of $8 USD per hour. Each annotator must pass benchmark tests and quality assurance checks before contributing to the project and a third reviewer (one of the authors) to verify annotation consistency and ensure pipeline quality.
Based on the Kappa scores reported in Table 3, most agreement levels fall within the moderate (0.41–0.60) to substantial (0.61–0.80) range (McHugh2012-hn). Assistant certainty and stance annotations show substantial and moderate agreement, respectively, for assistant-human conversations ( and ), and moderate agreement for assistant-chat data ( and ). For user-based annotations, user-chat data exhibited substantial agreement on user stance () and moderate agreement on implicit opinion (), with the highest average Kappa overall (). In contrast, user-human data showed only fair agreement on implicit opinion () and moderate agreement on user stance (), with the lowest overall average agreement ().